OpenStudy (anonymous):

Which function is an example of exponential decay? y = 0.3(2)x y = 4(8)x y = 5(0.6)x y = 1.4(3)x

OpenStudy (ybarrap):

Are any of these exponents?

OpenStudy (anonymous):

no @ybarrap

OpenStudy (the_fizicx99):

Ok, anything below 1 is decay. Anything above a 0.99 is growth :)

OpenStudy (the_fizicx99):

0.4 is the same as 0.40 or 40% btw

OpenStudy (anonymous):

???

OpenStudy (the_fizicx99):

K, um do you see those numbers in the parenthesis? y = 0.3(2) <--x y = 4(8)<--x y = 5(0.6)<--x y = 1.4(3)<--x Those are rates at which the x is then raise to, So suppose you're given to find f(6) in a random function like this, f(x) = 4(0.8)^x You carry the 6 inside the parenthesis, f(6) = 4(0.8^6) f(6) = 4(0.262144) f(6) = 1.048576 So for f(x) = 4(0.8)^x the 6th number in the sequence will be 1.048576 Now see that 0.8 ? Because its below 1 it'll be a decay. If you go increasing the number of x you'll see that it goes decreasing.

OpenStudy (the_fizicx99):

Usually if your told to make a funtion you use this, f(x)=P(1+r)^x You add 1 to r or rate, so 0.8 because 1.8

OpenStudy (anonymous):

can u help me on my hw???

OpenStudy (the_fizicx99):

._. did you figure out which one it is?

OpenStudy (anonymous):

ya so can u or no???

OpenStudy (the_fizicx99):

I can help you, but i can't give you the answers. Which one is it?

OpenStudy (anonymous):

ya I know

OpenStudy (the_fizicx99):

Well? A , B ,C or D ?

OpenStudy (anonymous):

b???

OpenStudy (the_fizicx99):

*shoots himself*

OpenStudy (the_fizicx99):

No U.U lol ANYTHING below 1 is decay. Anything ABOVE 0.99 is growth. Now which one is it :D

OpenStudy (anonymous):

okay gtg

OpenStudy (the_fizicx99):

._. tell me before you leave, a,b,c, or d D:

OpenStudy (the_fizicx99):

Btw, Just like its linear counterpart, exponential functions also have parameters, real world situation. f(x)=P(1+r)^x The P is the principal. This is the starting value, or y-intercept, when x = 0. The r is the rate of change. Remember that if the rate is a percentage, the r will be the decimal equivalent.